# Deflection angle in the strong deflection limit for static axisymmetric spacetimes: local curvature, matter fields, and quasinormal modes

### v2

@@ Acknowledgements.
The author gratefully acknowledges the useful discussions with Yohsuke Takamori. This work was supported in part by JSPS KAKENHI Grant Nos. JP22K03611, JP23KK0048, and JP24H00183 and by Gakushuin University.
@@ Appendix A Analysis of the $R^{\prime}_{\mathrm{m}}=0$ case
$\bar{b}$ $=\bar{a}\log\frac{4e^{2(\gamma_{\mathrm{m}}-\psi_{\mathrm{m}})}}{\bar{a}^{2}R_{\mathrm{m}}R^{\prime\prime}_{\mathrm{m}}}.$ (152)
@@
$\bar{b}=\bar{a}\log\left[\>\!\frac{16}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!2(E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}})-G_{(0)(0)}^{\mathrm{m}}-G_{(3)(3)}^{\mathrm{m}}+2G_{(2)(2)}^{\mathrm{m}}\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (153)
@@
$\bar{b}=\bar{a}\log\left[\>\!\frac{8}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-4\pi(T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}-2T_{(2)(2)}^{\mathrm{m}})\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (154)
@@
$\bar{b}=\bar{a}\log\left[\>\!\frac{8}{\bar{a}^{2}R_{\mathrm{m}}^{2}\left[\>\!-2(\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}}-\Phi_{11}^{\mathrm{m}}+\Phi_{02}^{\mathrm{m}})-\mathcal{R}^{\mathrm{m}}/4\>\!\right]}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (155)

### v4

@@ Acknowledgements.
The author gratefully acknowledges useful discussions with Yohsuke Takamori, Tomohiro Harada, and Akihito Katsumata. This work was supported in part by JSPS KAKENHI Grants No. JP22K03611, No. JP23KK0048, and No. JP24H00183 and by Gakushuin University.
DATA AVAILABILITY
No data were created or analyzed in this study.
@@ Appendix A Analysis of the $R^{\prime}_{\mathrm{m}}=0$ case
$\bar{b}$ $=-\bar{a}\log\left[\>\!\frac{\bar{a}^{2}}{4}R_{\mathrm{m}}R^{\prime\prime}_{\mathrm{m}}e^{2(\psi_{\mathrm{m}}-\gamma_{\mathrm{m}})}\>\!\right]+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (151)
@@
$\bar{b}=-\bar{a}\log\left\{\>\!\frac{\bar{a}^{2}}{16}R_{\mathrm{m}}^{2}\left[\>\!2(E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}})-G_{(0)(0)}^{\mathrm{m}}-G_{(3)(3)}^{\mathrm{m}}+2G_{(2)(2)}^{\mathrm{m}}\>\!\right]\>\!\right\}+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (152)
@@
$\bar{b}=-\bar{a}\log\left\{\>\!\frac{\bar{a}^{2}}{8}R_{\mathrm{m}}^{2}\left[\>\!E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-4\pi(T_{(0)(0)}^{\mathrm{m}}+T_{(3)(3)}^{\mathrm{m}}-2T_{(2)(2)}^{\mathrm{m}})\>\!\right]\>\!\right\}+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi,$ (153)
@@
$\bar{b}=-\bar{a}\log\left\{\>\!\frac{\bar{a}^{2}}{8}R_{\mathrm{m}}^{2}\left[\>\!-2(\Psi_{0}^{\mathrm{m}}+\Phi_{00}^{\mathrm{m}}-\Phi_{11}^{\mathrm{m}}+\Phi_{02}^{\mathrm{m}})-\frac{\mathcal{R}^{\mathrm{m}}}{4}\>\!\right]\>\!\right\}+I_{\mathrm{R}}(\rho_{\mathrm{m}})-\pi.$ (154)
Appendix B Outline of the derivation of $V^{\prime\prime}_{\mathrm{m}}$ in terms of curvature quantities
In this appendix, we provide a brief outline of how the coordinate derivatives in Eq. (12) are reorganized into the curvature quantities appearing in Eq. (55), using the tetrad components of the Einstein and Weyl tensors. Using the $\mathbb{Z}_{2}$ symmetry conditions (54) on the $\zeta=0$ plane, the tetrad components of the Einstein tensor and the Weyl tensor simplify considerably. In particular, the combination $G_{(0)(0)}^{\mathrm{m}}+G_{(3)(3)}^{\mathrm{m}}$ takes the form
$G_{(0)(0)}^{\mathrm{m}}+G_{(3)(3)}^{\mathrm{m}}=2e^{2\>\!(\psi_{\mathrm{m}}-\gamma_{\mathrm{m}})}\left[\>\!\psi_{\rho\rho}^{\mathrm{m}}+\psi_{\zeta\zeta}^{\mathrm{m}}+\frac{(W_{\rho}^{\mathrm{m}})^{2}}{2W_{\mathrm{m}}^{2}}-\frac{W_{\rho\rho}^{\mathrm{m}}+W_{\zeta\zeta}^{\mathrm{m}}}{2W_{\mathrm{m}}}\>\!\right],$ (155)
where we have used Eq. (11) to eliminate $\psi_{\rho}^{\mathrm{m}}$. Similarly, the difference of the tetrad components of the electric part of the Weyl tensor on the $\zeta=0$ plane reduces to the following expression:
$E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}=e^{2\>\!(\psi_{\mathrm{m}}-\gamma_{\mathrm{m}})}\left[\>\!\psi_{\zeta\zeta}^{\mathrm{m}}-\psi_{\rho\rho}^{\mathrm{m}}-\frac{(W_{\rho}^{\mathrm{m}})^{2}}{2W_{\mathrm{m}}^{2}}+\frac{W_{\rho\rho}^{\mathrm{m}}-W_{\zeta\zeta}^{\mathrm{m}}}{2W_{\mathrm{m}}}\>\!\right],$ (156)
where we have used Eq. (11). Then we have
$E_{(2)(2)}^{\mathrm{m}}-E_{(1)(1)}^{\mathrm{m}}-\frac{G_{(0)(0)}^{\mathrm{m}}+G_{(3)(3)}^{\mathrm{m}}}{2}=-\frac{e^{2\>\!(\psi_{\mathrm{m}}-\gamma_{\mathrm{m}})}}{W_{\mathrm{m}}^{2}}\left[\>\!2W_{\mathrm{m}}^{2}\psi_{\rho\rho}^{\mathrm{m}}+(W_{\rho}^{\mathrm{m}})^{2}-W_{\mathrm{m}}W_{\rho\rho}^{\mathrm{m}}\>\!\right].$ (157)
Substituting this result into Eq. (12) and using the definition of $R$ in Eq. (3), we immediately obtain $V^{\prime\prime}_{\mathrm{m}}$ in the form of Eq. (55).